A small remote-controlled car with mass 1.60 kg moves at a constant speed of v = 12.0 m/s in a track formed by a vertical circle inside a hollow metal cylinder that has a radius of 5.00 m, see Fig below. What is the magnitude of the normal force exerted on the car by the walls of the cylinder at

a) point A (bottom of the track)?

b) point B (top of the track)?
Answer #1

Given, m=1•6kg , V= 12 m/s , rasm B. Let assume Point A & B as, ta ca) Here, EF= n-mg = mue А mg R, nem +9.8 = 6.8N [X2+3) =